OpenStudy (anonymous):

(7x2-4x+22)+(3x^2+8x-89)

OpenStudy (anonymous):

7x2 ??? really 7 times 2, or the question is 7x^2 ???

OpenStudy (angela210793):

i think is 7x^2

OpenStudy (anonymous):

i think so too lol

OpenStudy (anonymous):

(7x^2-4x+22)+(3x^2+8x-89) = (7+3)x^2 + (-4+8)x + (22-89) = 10x^2+4x-67